OpenStudy (anonymous):

What is the 8th term of the geometric sequence -1,3,-9....?

OpenStudy (jhannybean):

what is your ratio?

OpenStudy (anonymous):

It doesn't say, would you like me to give you the choices?

OpenStudy (anonymous):

-59,049 -6,561 2,187 19,683

OpenStudy (jhannybean):

no,no,you have to FIND it. the radio is found by dividing the second term by the first.

OpenStudy (anonymous):

Oh, forgive me. I just started to learn this.

OpenStudy (anonymous):

So it would be -3

OpenStudy (anonymous):

\[a_n=a_1 \times r ^{n-1}\] a(n)=the nth term a(1)=the 1st term r=common ratio

OpenStudy (jhannybean):

Ohhh okay. Let me explain what a geometric sequence is. a Geoemtric sequence .... ok Kaus hahaha

OpenStudy (jhannybean):

so we found our r=-3 and we're to find our 8th term. if we follow our formula \[\large a_{n}=a_{1}r^{n-1}\] we'll be looking for our 8th number, with this, we find it by, \[\large a_{8}=(-1)(-3)^{8-1}\]\[\large a_{8}=(-1)(-3)^{7}=2187\]

OpenStudy (anonymous):

-6,561?

OpenStudy (anonymous):

ARUGH

OpenStudy (anonymous):

I need to study....

OpenStudy (jhannybean):

How did you get -6561? would you explain this to me? :P

OpenStudy (anonymous):

But thank you so much!

OpenStudy (anonymous):

Ill try another one on my own

OpenStudy (jhannybean):

Which calculus book are you using, by the way?Is it Stewart's?

OpenStudy (anonymous):

I dont have a textbook, I work and take classes online.

OpenStudy (jhannybean):

Yeah, first find your ratio, your "a1" or first number, and then plug it into the equation to find the nth term, whatever you're looking for.

OpenStudy (jhannybean):

Ohh nevermind, I was going to direct you to the chapter for further study, haha.

OpenStudy (anonymous):

SO this sequence is a1 = 1024 and a3=64 and I got 0.0625 Hoping thats correct? @Jhannybean

OpenStudy (jhannybean):

which sequence? geometric? new problem?

OpenStudy (anonymous):

New problem and yes geometric, but i already check and it was correct (:

OpenStudy (jhannybean):

Yay!! you're learning this stuff already!